OpenStudy (anonymous):

Im solving systems: how do i solve 5x+3y=-9 3x-5y=-9.

OpenStudy (alexwee123):

use substitution or elimination or graph or matrices :/

OpenStudy (alexwee123):

your choice XD

OpenStudy (anonymous):

Okay,thank you!

OpenStudy (anonymous):

x = -36/17 y=9/17